OpenStudy (anonymous):

Find a set of the vectors of the reciprocal to the set: 3i - j + 2k , 2i + j - k and i - 2j + k

OpenStudy (anonymous):

Given three non coplanar vectors a, b, c, the reciprocal vectors A, B, C are: A = (b x c)/w B = (c x a)/w C = (a x b)/w Where w = a . (b x c) is the volume of the parallelepiped fromed by a, b, and c. Note: a x c means the cross product of the vectors a and c. Now just plug the numbers in.

OpenStudy (anonymous):

like i dont understand how we going to work with the fraction

OpenStudy (turingtest):

do you know how to find the volume of the parallelepiped Mr._To mentioned ? that is W in your formula...\[W=\vec a\cdot(\vec b\times\vec c)\]

OpenStudy (anonymous):

guys please help...

OpenStudy (anonymous):

what i know is that a . ( b x c ) is not equal to zero

OpenStudy (turingtest):

well first you need\[\vec b\times\vec c\]do you know how to take the cross-product of two vectors?

OpenStudy (anonymous):

yeah...

OpenStudy (turingtest):

then that is step one

OpenStudy (anonymous):

ohk let me work the answer and you tell if am wrong

OpenStudy (turingtest):

let's say\[\vec a=3\hat i-\hat j+2\hat k\]\[\vec b=2\hat i+\hat j-\hat k\]\[\vec c=\hat i-2\hat j+\hat k\]so you tell me what \[\vec b\times\vec c\]is

OpenStudy (anonymous):

-i - 3j - 5k

OpenStudy (turingtest):

yeah, that's what I got :) so what's\[\vec a\cdot(\vec b\times\vec c)\]?

OpenStudy (anonymous):

10

OpenStudy (turingtest):

-10 ain't it?

OpenStudy (anonymous):

oh yeah -10

OpenStudy (turingtest):

so now we know W=-10 you know how to take the cross-product of two vectors, so you should be able to use the formula @Mr._To provided

OpenStudy (turingtest):

\[\vec A={\vec b\times\vec c\over W}\]\[\vec B={\vec c\times\vec a\over W}\]\[\vec C={\vec a\times\vec b\over W}\]

OpenStudy (anonymous):

so in these quiz am looking for A B C

OpenStudy (turingtest):

yup

OpenStudy (anonymous):

So now cos the ( b x c ) is going to give me answer that got directions what am i suppose to do?

OpenStudy (turingtest):

well \(\vec b\times\vec c\) is a vector, which is what we want so now we just scalar divide that vector by W

OpenStudy (anonymous):

which is the scalar?

OpenStudy (turingtest):

W scalar is a number without direction; basically the opposite of a vector

OpenStudy (turingtest):

scalar multiplication on a vector is straightforward for each component

OpenStudy (turingtest):

\[c\vec v=c\langle x,y,z\rangle=\langle cx,cy,cz\rangle\]

OpenStudy (turingtest):

division works the same way of course

OpenStudy (anonymous):

so i take the W to divide the vector?

OpenStudy (turingtest):

yes, you are scalar multiplying each vector that you get from the cross-product by 1/W (which is the same thing as dividing by W....) to get the inverse

OpenStudy (turingtest):

we already know \(\vec b\times\vec c\) from our earlier work so now just scalar divide that by W

OpenStudy (turingtest):

.... in order to find \(\vec A\)

OpenStudy (anonymous):

lets say for example i go i - 2j + 5k i take w x i , w x (-2 )...

OpenStudy (turingtest):

well the formula Mr._To has given is to divide by W, not multiply but you could write it like that

OpenStudy (anonymous):

ohk... so the answer suppose to be a vector or scalar?

OpenStudy (turingtest):

better to be consistent in your notation; either\[\vec A=\frac1W(x\hat i+y\hat j+z\hat k)=\frac xW\hat i+\frac yW\hat j+\frac zW\hat k\]or\[\vec A=\frac1W\langle x,y,z\rangle=\langle \frac xW,\frac yW,\frac zW\rangle\]as you can see our answers are vectors

OpenStudy (anonymous):

ohkay... i will try it out

OpenStudy (anonymous):

Turing Test, great instructions!

OpenStudy (turingtest):

Just following your lead Mr._To, that formula was news to me!